# Association of the IceCube neutrinos with CAZ blazar light curves

### v1

@@
Over the past decade, the IceCube Neutrino Observatory has detected a few hundreds of high-energy (HE) neutrinos from cosmic sources. Despite numerous studies searching for their origin, it is still not known which source populations emit them. A few confident individual associations exist with active galactic nuclei (AGN), mostly with blazars which are jetted AGN whose jet points in our direction. Nonetheless, on a population level, blazar-neutrino correlation strengths are rather weak. This could mean that blazars as a population do not emit HE neutrinos, or that the detection power of the tests is insufficient due to the strong atmospheric neutrino background. By assuming an increase in HE neutrino emission during major blazar flares, in our previous studies we leveraged the arrival time of the neutrinos to boost the detection power. In this paper we utilize the same principle while substantially increasing the number of blazars. We search for the spatio-temporal correlation of 356 IceCube HE neutrinos with major optical flares of 3225 radio- and 3814 -ray-selected blazars. We find that, despite the increase in data size, the number of confident spatio-temporal associations remains low and the overall correlation strengths weak. Two individual associations drive our strongest and the only 2 post-trial spatio-temporal correlation, occurring with the BL Lac objects of the radio-selected blazar sample. We estimate that 8% of the detected cosmic neutrinos were emitted by blazars during major optical flares. As a complementary analysis, we compare the synchrotron peak frequency, redshift, Doppler factor, X-ray brightness, and optical variability of spatially neutrino-associated blazars to those of the general blazar population. We find that spatially neutrino-associated blazars of the tested samples have higher than average Doppler factor and X-ray brightness.
@@
As of now, hundreds of high-energy (TeV-PeV) neutrinos from extraterrestrial sources have been detected in excess to the strong atmospheric neutrino background by various neutrino observatories, such as IceCube (e.g., IceCube Collaboration et al. 2021), ANTARES (e.g., Albert et al. 2021), and KM3NeT (e.g., KM3NeT Collaboration et al. 2025b), and more are expected to be detected by the next generation or future observatories, such as Baikal-GVD (e.g., Baikal-GVD Collaboration et al. 2018), RNO-G (e.g., Aguilar et al. 2021), GRAND (e.g., Fang et al. 2017), and P-ONE (e.g., Malecki 2024). Despite hundreds of such detections already, the cosmic sources that produce high-energy neutrinos are still largely unknown (e.g., Abbasi et al. 2021b; Troitsky 2021).
@@
One of the main candidates for emitting high-energy (TeV–PeV) neutrinos (hereafter “neutrinos”) are active galactic nuclei (AGN; e.g., Mannheim & Biermann 1989; Mücke & Protheroe 2001; Hooper & Plant 2023). Powered by supermassive black holes, almost all AGN display extreme nonthermal emission and 10% possess highly collimated jets of relativistic plasma. AGN jets could accelerate hadrons to energies high enough to produce neutrinos (e.g., Mannheim 1993). In AGN whose jets are viewed at small angles (; i.e., blazars), all electromagnetic and any potential neutrino emission from the jet becomes strongly Doppler boosted. This makes blazars some of the most luminous objects in the universe and prime targets for a population-based neutrino counterpart search. For recent reviews on blazars, see, for example, Böttcher (2019), Blandford et al. (2019), and Hovatta & Lindfors (2019).
@@
While the hottest neutrino spot in the sky is NGC 1068 (a weakly jetted Seyfert-II AGN, e.g., Muxlow et al. 1996; Padovani et al. 2024b) at a significance of 4.2 (IceCube Collaboration et al. 2022), the only other few sources confidently associated with neutrinos have been blazars (e.g., TXS 0506+056, IceCube Collaboration et al. 2018; PKS 1424+240, and GB6 J1542+6129, Abbasi et al. 2021a). As such, many studies have focused on the connection of neutrinos with individual blazars (e.g., Krauß et al. 2014; Kadler et al. 2016; Padovani et al. 2019; Righi et al. 2019; Franckowiak et al. 2020; Rodrigues et al. 2021; Padovani et al. 2022; Acharyya et al. 2023; Kun et al. 2024; Blinov & Novikova 2025; Omeliukh et al. 2025; KM3NeT Collaboration et al. 2025a; Kovalev et al. 2025) and blazar populations (e.g., Aartsen et al. 2020; Plavin et al. 2020; Giommi et al. 2020; Plavin et al. 2021; Smith et al. 2021; Hovatta et al. 2021; Zhou et al. 2021; Bartos et al. 2021; Kun et al. 2022; Buson et al. 2022; Plavin et al. 2023; Novikova et al. 2023; Buson et al. 2023; Bellenghi et al. 2023; Suray & Troitsky 2024; Albert et al. 2024; Lu et al. 2024; Plavin et al. 2024; Kouch et al. 2024a; Abbasi et al. 2024b, 2025). Some of these population-based analyses hint toward a potential global connection between blazars and neutrinos, albeit rather inconclusively.
@@
There is evidence that the most energetic neutrinos are not coming from continuously neutrino-emitting sources, but possibly from a source population that mainly and rarely emits the highest energy neutrinos (e.g., Abbasi et al. 2024a). Blazars are suspected to be such a source population (e.g., Padovani et al. 2024a). Yet, a confident population-based correlation between blazars and these highest energy neutrinos has not been established. This may be the result of: (1) only a few blazars emitting such neutrinos; or (2) inadequate detection power due to suboptimal test statistics and observational limitations (i.e., low counts, low signalness111Signalness is an estimate for the probability of a neutrino event being of cosmic origin., and poor spatial resolution of the reconstructed neutrino events; e.g., Liodakis et al. 2022a; Kouch et al. 2025b). To understand whether (1) is the reason and to confidently identify potential neutrino-emitting blazars, it is vital to diminish the effect of (2), that is, to boost detection power.
@@
In our previous works (Hovatta et al. 2021 and Kouch et al. 2024a, hereafter “H21” and “K24”, respectively), we attempted to boost the detection power of our correlation analysis by leveraging the precise arrival time of neutrinos (e.g., Abbasi et al. 2024b). Since blazar emission is highly variable, with quiescent and flaring periods generally corresponding to lower and higher jet activity, respectively (e.g., Lähteenmäki & Valtaoja 2003; León-Tavares et al. 2011), we assumed that neutrino production would be enhanced when the jet is more active during major electromagnetic flares (e.g., Oikonomou et al. 2019; Kreter et al. 2020; Stathopoulos et al. 2022). Hence, we searched for temporal associations between individual high-energy neutrino arrival times and major, long-term flares experienced by spatially associated blazars in the radio (in both H21 and K24) and optical bands (in K24). We obtained post-trial spatio-temporal correlations of 2 and 2.17 in H21 and K24, respectively.
@@
In this paper, while using a similar neutrino sample as K24 (i.e., IceCat1+, but updated to include 2021–2023 neutrino events; see Sect. 2.1), we test the connection between neutrinos and blazars using short-term optical variability and flares. Moreover, we now enhance the detection power of our spatio-temporal analysis even further by: (1) using a larger sample of blazar light curves which we compiled in our companion paper Kouch et al. (2025a), hereafter “K25-Cat” (e.g., see our simulations in Liodakis et al. 2022a which suggested that having more sources should result in a more significant signal if a spatio-temporal blazar-neutrino connection exists); and (2) using the most optimal test strategy as determined via our simulations in Kouch et al. (2025b), hereafter “K25-Sim”. We note that the most optimal test strategy was also utilized in K24.
@@ 2.1 High-energy neutrino events
In this study we use an updated version of the IceCat1+ neutrino sample which we compiled in K24. IceCat1+ consists of 267 neutrinos from the IceCube collaboration’s IceCat-1 (Abbasi et al. 2023b), which is the first catalog of muon-track high-energy neutrino events selected based on real-time alert criteria (recorded between May 2011 and December 2020). IceCat1+ additionally contains 16 muon-track events from another IceCube collaboration study (Abbasi et al. 2022) where events recorded between May 2009 (when the telescope was still in partial configuration) and December 2018 were selected based on non-real-time criteria. In this paper, we additionally include 73 events between January 2021 and October 2023 which were added to IceCat-1 after its initial publication (v4.0222https://dataverse.harvard.edu/file.xhtml?fileId=7502710). Therefore, this “updated IceCat1+” comprises 356 events in total and is given as an electronic table in this paper. This table provides the arrival time, likely sky localization, kinetic energy estimate, and signalness () of the 356 events.
While the arrival times are known precisely (on the order of nanoseconds), the sky localizations are not. Using the reported asymmetric right ascension (RA) and declination (Dec.) errors as one-sided 2 Gaussian error estimates, one can trace an ellipsoid corresponding to the 90%-likelihood error region of each event in the sky (see Fig. 1 of K25-Sim). The area of such an error region () is calculated as follows:
@@
Notably, the median signalness for the updated IceCat1+ is , implying that more than half of its neutrinos are of atmospheric origin. These atmospheric neutrinos act as noise when searching for a global blazar-neutrino association. In Sect. 3 we describe a test strategy which involves weighing down neutrinos with large and small to optimally boost the detection power of the blazar-neutrino search.
@@ 2.2 Blazars, their light curves, and their flares
In this study we use two blazar-dominated AGN samples: (1) a statistically complete subsample of the Radio Fundamental Catalog containing 3225 sources (hereafter denoted as RFC†; see Sect. 2.2.1); and (2) the Fourth Catalog of Active Galactic Nuclei present in the Large Array Telescope (4LAC) containing 3814 sources (see Sect. 2.2.2). Both of these are substantially larger than the blazar samples we used in our previous studies. In H21 we used several samples with a combined count of 1795, whereas in K24 we only focused on the CGRaBS sample (the best from H21) with a count of 1157.
@@
The RFC† and 4LAC samples are both contained in the CAZ catalog, constituting a total of 7918 sources, which we compiled in K25-Cat. There, we also tabulated, when available, their physical parameters such as source type, synchrotron peak frequency (), redshift (), radio variability Doppler factor (), and average X-ray flux density (). We provide their X-ray light curves and spectra in another companion paper Paggi et al. (2025). In the CAZ catalog, the source types are: flat spectrum radio quasar (FSRQ; denoted as “Q”), BL Lac object (BLL; denoted as “B”), host-galaxy dominated BLL (denoted as “G”), blazar of unknown type (denoted as “U”), or non-blazar AGN (denoted as “A”). We additionally divide the sources into classes based on their following the low-, intermediate-, high-, and extremely high-synchrotron peaked scheme (abbreviated as LSP, ISP, HSP, and EHSP, respectively). LSPs have Hz, ISPs Hz, HSPs Hz, and EHSPs Hz. We note that FSRQs are generally LSPs, but BLLs can range from LSPs to EHSPs and, thus, are further divided into: LBLs, IBLs, HBLs, and EHBLs. In Table 1 we summarize the number and fraction of available parameters for the RFC† and 4LAC samples.
@@
In K25-Cat we additionally constructed decade-long, high-cadence optical light curves for most of the 7918 CAZ sources using the Catalina Real-time Transient Survey (CRTS), the Asteroid Terrestrial-impact Last Alert System (ATLAS), and the Zwicky Transient Facility (ZTF) all-sky light curves, which are collectively referred to as the CAZ light curves. We also identified their variability (see Sect. 2.2.3), periods of enhanced emission (see Sect. 2.2.4), and prominent flaring periods (see Sect. 2.2.5 and 2.2.6). These light curve and variability information along with some of the physical parameters of the blazars are used in the spatio-temporal analysis of this study. Crucially, only sources with CAZ light curves deemed variable are considered in the spatio-temporal analysis. However, in Sect. 4.7, we perform a complementary, spatial-only analysis using all blazars regardless of their variability.
@@ 2.2.1 RFC† sample
The Radio Fundamental Catalog (RFC444https://astrogeo.org/sol/rfc/) contains tens of thousands of precisely localized compact radio sources detected via very long baseline interferometry (VLBI). A subsample of it, limited to mJy, can be considered statistically complete as a result of observations by the VCS (Very Long Baseline Array calibrator surveys; e.g., Kovalev et al. 2007; Gordon et al. 2016) and LBA (Australian Long Baseline Array; e.g., Petrov et al. 2019). Since flux-limited VLBI samples are generally blazar-dominated (e.g., Readhead et al. 1978), we began compiling the CAZ catalog using this subsample of RFC555The rfc_2023d version of RFC was used in K25-Cat.. The sources in the CAZ catalog were checked to be: (1) AGN; and (2) unique within 0.001∘ (see K25-Cat for justification).
Of the 7918 CAZ sources, 3225: (1) are in RFC; (2) are categorized as AGN or candidates; and (3) fulfill mJy. These 3225 sources form the statistically complete RFC sample (i.e., RFC†), whose parameter availability is summarized in Table 1 column (1). We note that RFC† is dominated by FSRQs and LBLs, as expected from a radio flux-limited sample. There are 1159 sources in RFC† which are also found in the 4LAC sample.
@@
4LAC is the second blazar-dominated sample we use in this study, which is also contained in its entirety in the CAZ catalog (see K25-Cat). We adopt all 3814 low- and high-latitude sources in the third 4LAC data-release (4LAC-DR3666https://fermi.gsfc.nasa.gov/ssc/data/access/lat/4LACDR3/) from August 2022, which is based on 12 years of data from the Fermi Large Array Telescope (LAT; Atwood et al. 2009). Crucially, we use the radio counterpart coordinates of the 4LAC sources instead of the less reliably constrained -ray coordinates (Ajello et al. 2022).

### v2

@@
The IceCube Neutrino Observatory has detected several hundred high-energy neutrinos from cosmic sources. Despite numerous studies searching for their origin, it is still not known which sources emit them. A few likely individual associations exist with active galactic nuclei (AGNs), mostly comprising blazars (AGNs with jets pointed toward Earth). Nonetheless, on a population level, blazar-neutrino correlation strengths are rather weak. This could mean that blazars as a population do not emit high-energy neutrinos or that the detection power of the tests is insufficient due to the strong atmospheric neutrino background. By assuming an increase in high-energy neutrino emission during major blazar flares, in our previous studies, we leveraged the neutrino arrival time to boost the detection power. Here, we utilize the same principle, while substantially increasing the number of blazars in the sample. We searched for the spatiotemporal correlation of 356 IceCube high-energy neutrinos with major optical flares of 3225 radio- and 3814 -ray-selected blazars. We found that despite the increase in data size, the number of likely spatiotemporal associations remained low and the overall correlation strengths weak. Two individual associations were shown to drive our strongest correlation, namely, the only 2 post-trial spatiotemporal correlation, occurring with the BL Lac objects of the radio-selected blazar sample. We estimated that 8% of the detected cosmic neutrinos were emitted by blazars during major optical flares. As a complementary analysis, we compared the synchrotron peak frequency, redshift, Doppler factor, X-ray brightness, and optical variability of spatially neutrino-associated blazars to those of the general blazar population. We found that spatially neutrino-associated blazars have a Doppler factor and X-ray brightness that are higher than average.
@@
At present, hundreds of high-energy (TeV-PeV) neutrinos from extraterrestrial sources have been detected and found to be in excess of the strong atmospheric neutrino background by various neutrino observatories, such as IceCube (e.g., IceCube Collaboration et al. 2021), ANTARES (e.g., Albert et al. 2021), and KM3NeT (e.g., KM3NeT Collaboration et al. 2025b). Further detections are expected by the next generation or future observatories, such as Baikal-GVD (e.g., Baikal-GVD Collaboration et al. 2018), RNO-G (e.g., Aguilar et al. 2021), GRAND (e.g., Fang et al. 2017), P-ONE (e.g., Malecki 2024), TRIDENT (e.g., Ye et al. 2023), and HUNT (e.g., Huang et al. 2024). Despite the hundreds of detections that are available, the cosmic sources that produce high-energy neutrinos are still largely unknown (e.g., Abbasi et al. 2021b; Troitsky 2021).
@@
One of the main candidates for emitting high-energy (TeV–PeV) neutrinos (hereafter, neutrinos) are active galactic nuclei (AGNs; e.g., Mannheim & Biermann 1989; Mücke & Protheroe 2001; Hooper & Plant 2023). Powered by supermassive black holes, almost all AGNs display extreme nonthermal emission and 10% possess highly collimated jets of relativistic plasma. AGN jets have the ability to accelerate hadrons to high-enough energies which could interact with photons to produce neutrinos (e.g., Mannheim 1993). In AGNs whose jets are viewed at small angles (; i.e., blazars), all electromagnetic and any potential neutrino emission from the jet becomes strongly Doppler boosted. This makes blazars some of the most luminous objects in the universe and prime targets for a population-based neutrino counterpart search. For recent reviews on blazars, see, for example, Böttcher (2019), Blandford et al. (2019), and Hovatta & Lindfors (2019).
@@
While the hottest neutrino spot in the sky is NGC 1068 (a weakly jetted Seyfert-II AGN, e.g., Muxlow et al. 1996; Padovani et al. 2024b) at a significance of 4.2 (IceCube Collaboration et al. 2022), the only other few sources potentially associated with neutrinos have been blazars (e.g., TXS 0506+056, IceCube Collaboration et al. 2018; PKS 1424+240, and GB6 J1542+6129, Abbasi et al. 2021a). As such, many studies have focused on the connection of neutrinos with individual blazars (e.g., Krauß et al. 2014; Kadler et al. 2016; Padovani et al. 2019; Righi et al. 2019; Franckowiak et al. 2020; Rodrigues et al. 2021; Padovani et al. 2022; Acharyya et al. 2023; Kun et al. 2024; Blinov & Novikova 2025; Omeliukh et al. 2025; KM3NeT Collaboration et al. 2025a; Kovalev et al. 2025) and blazar populations (e.g., Aartsen et al. 2020; Plavin et al. 2020; Giommi et al. 2020; Plavin et al. 2021; Smith et al. 2021; Hovatta et al. 2021; Zhou et al. 2021; Bartos et al. 2021; Kun et al. 2022; Buson et al. 2022; Plavin et al. 2023; Novikova et al. 2023; Buson et al. 2023; Bellenghi et al. 2023; Suray & Troitsky 2024; Albert et al. 2024; Plavin et al. 2024; Kouch et al. 2024a; Abbasi et al. 2024b, 2025; Lu et al. 2025). Some of these population-based analyses hint toward a potential global connection between blazars and neutrinos, albeit rather inconclusively.
@@
There is evidence that the most energetic neutrinos are not coming from continuously neutrino-emitting sources, but possibly from a source population that mainly and rarely emits the highest energy neutrinos (e.g., Abbasi et al. 2024a). Blazars are suspected to be such a source population (e.g., Padovani et al. 2024a). Yet, a confident population-based correlation between blazars and these highest energy neutrinos has not been established. This may be the result of: (1) only a few blazars emitting such neutrinos or (2) inadequate detection power due to suboptimal test statistics and observational limitations (i.e., low counts, low signalness,111Signalness is an estimate for the probability of a neutrino event being of cosmic origin. and poor spatial resolution of the reconstructed neutrino events; e.g., Liodakis et al. 2022a; Kouch et al. 2025a). To understand whether (1) is the reason and to confidently identify potential neutrino-emitting blazars, it is vital to diminish the effect of (2) by boosting the detection power.
@@
In our previous works (Hovatta et al. 2021 and Kouch et al. 2024a, hereafter “H21” and “K24,” respectively), we attempted to boost the detection power of our correlation analysis by leveraging the precise arrival time of neutrinos (e.g., Abbasi et al. 2024b). Since blazar emission is highly variable, with quiescent and flaring periods generally corresponding to lower and higher jet activity, respectively (e.g., Lähteenmäki & Valtaoja 2003; León-Tavares et al. 2011), we assumed that neutrino production would be enhanced when the jet is more active during major electromagnetic flares (e.g., Oikonomou et al. 2019; Kreter et al. 2020; Stathopoulos et al. 2022). Via this phenomenological assumption, we searched for temporal associations between individual high-energy neutrino arrival times and major, long-term flares experienced by spatially associated blazars in the radio (in both H21 and K24) and optical bands (in K24). We obtained post-trial spatiotemporal correlations of 2 and 2.17 in H21 and K24, respectively.
@@
In this paper, we use a similar neutrino sample as K24 (i.e., IceCat1+, but updated to include 2021–2023 neutrino events; see Sect. 2.1) and rely on the same phenomenological assumption as above (i.e., in blazars, major neutrino flares coincide with major electromagnetic flares). However, unlike the approach taken in H21 and K24, here we search for quasi-simultaneous coincidences between neutrino arrival times and major, short-term optical flares (some as short as a few days in duration). We focus on variability in the optical band because: (1) most blazars are variable in the optical band, with optical flaring being often correlated with flaring in higher energy bands (e.g., -rays; Liodakis et al. 2019; de Jaeger et al. 2023) and (2) the optical band offers the largest number of well-sampled light curves, which is a necessary ingredient for performing temporal associations, owing to recent advances in all-sky surveys. Crucially, we can now enhance the detection power of our spatiotemporal analysis even further by: (1) using the largest sample of blazar optical light curves to date, compiled in our companion paper Kouch et al. (2025b), hereafter referred to as K26 (e.g., see our simulations in Liodakis et al. 2022a which suggested that having more sources should result in a more significant signal if a spatiotemporal blazar-neutrino connection exists) and (2) using the most optimal test strategy as determined via our simulations in Kouch et al. (2025a), hereafter referred to as “K25”. We note that the most optimal test strategy was also utilized in K24.
@@ 2.1 High-energy neutrino events
In this study, we use an updated version of the IceCat1+ neutrino sample, which we previously compiled in K24. IceCat1+ consists of 267 neutrinos from the IceCube collaboration’s IceCat-1 (Abbasi et al. 2023b), which is the first catalog of muon-track high-energy neutrino events selected based on real-time alert criteria (recorded between May 2011 and December 2020). IceCat1+ additionally contains 16 muon-track events from another IceCube collaboration study (Abbasi et al. 2022), where events recorded between May 2009 (when the telescope was still in partial configuration) and December 2018 were selected based on non-real-time criteria. In this paper, we additionally include 73 events between January 2021 and October 2023, which were added to IceCat-1 after its initial publication (v4.0222https://dataverse.harvard.edu/file.xhtml?fileId=7502710). Therefore, this “updated IceCat1+” comprises 356 events in total and is given as an electronic table in this paper. This table provides the arrival time, likely sky localization, kinetic energy estimate, and signalness () of the 356 events.
While the arrival times are known precisely (on the order of nanoseconds), the sky localizations are not. Using the reported asymmetric right ascension (RA) and declination (Dec.) errors as one-sided 2 Gaussian error estimates, we can trace an ellipsoid corresponding to the 90%-likelihood error region of each event in the sky (see Fig. 1 of K25). The area of such an error region () is calculated as follows:
@@
Notably, the median signalness for the updated IceCat1+ is , implying that more than half of its neutrinos are of atmospheric origin. These atmospheric neutrinos act as noise when searching for a global blazar-neutrino association. In Sect. 3, we describe a test strategy that involves weighing down neutrinos with large and small to optimally boost the detection power of the blazar-neutrino search.
@@ 2.2 Blazars, their light curves, and their flares
In this study, we used two blazar-dominated AGN samples: (1) a statistically complete subsample of the Radio Fundamental Catalog containing 3225 sources (hereafter denoted as RFC†; see Sect. 2.2.1) and (2) the Fourth Catalog of Active Galactic Nuclei present in the Large Array Telescope (4LAC) containing 3814 sources (see Sect. 2.2.2). Both of these are substantially larger than the blazar samples we used in our previous studies. In H21 we used several samples with a combined count of 1795; whereas in K24, we only focused on the CGRaBS sample (the best from H21) with a count of 1157.
@@
In K26, we compiled the largest blazar catalog to date, consisting of 7918 unique sources from the RFC†, 4LAC, third catalog of HSP (3HSP), Candidate Gamma-Ray Blazar Survey (CGRaBS), and Fifth Roma-BZCAT Multifrequency Catalogue of Blazars (5BZC) catalogs. Notably, the RFC† and 4LAC samples are both fully contained in the CAZ catalog. Together they constitute 5880 sources, making up the majority (74%) of the CAZ catalog. In this study, we use these two subsamples of the CAZ catalog instead of its entirety in order to investigate whether radio- or -ray-bright blazars in particular would be more likely to be neutrino emitters. In K26, we also tabulated, when available, the physical parameters of the CAZ sources, such as source type, synchrotron peak frequency (), redshift (), radio variability Doppler factor (), and average X-ray flux density (). We provide their X-ray light curves and spectra in another companion paper Paggi et al. (in prep.). In the CAZ catalog, the source types are: flat spectrum radio quasar (FSRQ; denoted as Q), BL Lac object (BLL; denoted as B), host-galaxy dominated BLL (denoted as G), blazar of unknown type (denoted as U), or non-blazar AGN (denoted as A). We additionally divide the sources into classes based on their following the low-, intermediate-, high-, and extremely high-synchrotron peaked scheme (abbreviated as LSP, ISP, HSP, and EHSP, respectively). LSPs have Hz, ISPs Hz, HSPs Hz, and EHSPs Hz. We note that FSRQs are generally LSPs, but BLLs can range from LSPs to EHSPs and are, thus, further divided into: LBLs, IBLs, HBLs, and EHBLs. In Table 1, we summarize the number and fraction of available parameters for the RFC† and 4LAC samples.
@@
In K26, we additionally constructed decade-long, high-cadence optical light curves for most of the 7918 CAZ sources using the Catalina Real-time Transient Survey (CRTS), the Asteroid Terrestrial-impact Last Alert System (ATLAS), and the Zwicky Transient Facility (ZTF) all-sky light curves, which are collectively referred to as the CAZ light curves. Their absolute flux density levels are not always indicative of the flux density of the jet. The host-galaxy and non-jet components of the blazar can contaminate the observed flux densities. Likewise, nearby bright sources can cause flux density contamination as a result of the unsupervised data reduction procedures of the all-sky surveys (e.g., automatic forced-photometry). Therefore, we used the relative changes in flux density rather than the absolute values to trace the intrinsic changes in the activity level of a blazar jet. As such, we identified the variability (see Sect. 2.2.3), periods of enhanced emission (see Sect. 2.2.4), and prominent flaring periods (see Sect. 2.2.5 and 2.2.6) of the CAZ light curves. These light-curve and variability data, along with some of the physical parameters of the blazars, were used in the spatiotemporal analysis of this study. Crucially, only sources with CAZ light curves deemed variable were considered in the spatiotemporal analysis. However, in Sect. 4.7, we describe our complementary, spatial-only analysis using all blazars regardless of their variability.
@@ 2.2.1 RFC† sample
The Radio Fundamental Catalog (RFC444https://astrogeo.org/sol/rfc/) contains tens of thousands of precisely localized compact radio sources detected via very long baseline interferometry (VLBI). A subsample of it, limited to mJy, can be considered statistically complete as a result of observations by the VCS (Very Long Baseline Array calibrator surveys; e.g., Kovalev et al. 2007; Gordon et al. 2016) and Australian LBA (Australian Long Baseline Array; e.g., Petrov et al. 2019). Since flux-limited VLBI samples are generally blazar-dominated (e.g., Readhead et al. 1978), we began compiling the CAZ catalog using this subsample of RFC555The rfc_2023d version of RFC was used in K26.. The sources in the CAZ catalog were checked to be: (1) AGN; and (2) unique within 0.001∘ (see K26 for justification).
Of the 7918 CAZ sources, 3225: (1) are in RFC; (2) are categorized as AGNs or candidates; and (3) fulfill mJy. These 3225 sources form the statistically complete RFC sample (i.e., RFC†), whose parameter availability is summarized in Table 1 column (1). We note that RFC† is dominated by FSRQs and LBLs, as expected from a radio flux-limited sample. There are 1159 sources in RFC† that are also found in the 4LAC sample.
@@
4LAC is the second blazar-dominated sample we use in this study, which is also contained in its entirety in the CAZ catalog (see K26). We adopted all 3814 low- and high-latitude sources in the third 4LAC data-release (4LAC-DR3666https://fermi.gsfc.nasa.gov/ssc/data/access/lat/4LACDR3/) from August 2022, which is based on 12 years of data from the Fermi Large Array Telescope (LAT; Atwood et al. 2009). Crucially, we used the radio counterpart coordinates of the 4LAC sources instead of the less reliably constrained -ray coordinates (Ajello et al. 2022).